OpenStudy (anonymous):

An investment of $7,650 earns interest at the rate of 5% and is compounded quarterly. What is the accumulated value of the investment at the end of 9 years? a. $11,964.17 b. $11,867.66 c. $9,338.56 d. $8,554.94

OpenStudy (anonymous):

its $11,964.17 bc \[7650 \times (1 + \frac{ 0.05 }{ 4 }) 4 \times 9 = 11964.17\]

OpenStudy (anonymous):

thank you @plohrr

OpenStudy (anonymous):

mhm c;